OpenStudy (anonymous):

Essay; show all work. A simple dartboard has three areas… the main board has a radius of 10 inches, there is a circle with a radius of 7 inches, and the bullseye has a radius of 3 inches. What is the probability Round to the nearest thousandths A=pi r^2 please show work!

OpenStudy (anonymous):

it is the ratio of the areas.

OpenStudy (anonymous):

but you are missing something important. What is the probability...

OpenStudy (anonymous):

probability of what?

OpenStudy (anonymous):

oh i forgot to post the picture!

OpenStudy (anonymous):

picture doesn't help probability of what?

OpenStudy (anonymous):

if you copied and pasted the question it did not come out correctly. missing a piece

OpenStudy (anonymous):

What is the probability of a random dart landing inside the bullseye?

OpenStudy (anonymous):

ok. now we can do it. "random" means you don't aim, so it can be anywhere. take the ratio of the areas.

OpenStudy (anonymous):

area of bulls eye is \[9\pi\] area of dartboard is \[100\pi\] ratio is \[\frac{9}{100}=.09\] (the pi cancels)